Q.6 Three year bonds are issued at face value of SR 100,000 on Jan. 1, 2007, and a stated interest rate of 8%. Calculate the issue price of the bonds assuming a market interest rate of 6%. The present value of SR 1 is .83962 at 6% after 3 years and the present value of an ordinary annuity of SR 1 is 2.67301 at 6% after 3 years.

Q.7. Three year 8% bonds of SR 100,000 issued on Jan. 1, 2007, are recalled at 105 on Dec. 31, 2008. Expenses of recall are SR 2,000. Market interest on issue date was 8%.

Q8. On Jan. 1, 2010, General Bell Corporation issued at 97%, bonds with a par value of SR 800,000 due in 20 years. It incurred bond issue cost totaling SR 16,000. Eight years after the issue date, General Bells calls the entire issue at 101% and cancels it. Compute the loss on redemption (extinguishment).   

Q.9. On Jan. 1, 2011, STC retired SR 500,000 of bonds at 99%. At the time of retirement, the unamortized premium was SR 15,000 and unamortized bond issue costs were SR 5,250. Prepare journal entries of reacquisition of bonds in the books of STC.      
Answer #1

Solution to Question 6 :

Calculation of Bond Price :

Price = 100,000*8%*( Annuity - Market Rate for Maturity period) + 100,000 ( PV factor Market rate for Maturity Period)

Price = 100,000*8%*2.67301 + 100,000*0.83962

Price = 8000*2.67301 + 83,962

Price = 21,384.08 + 83,962

Price = 105,346.08

The price of the bond is more than the face value on the account of decrease in the rates in bond market. This makes the bond theoretically more attractive as it is paying high interest as compared to the other options available in the market.

Note : The solution to one Question can be provided only at a time. Also, Q7 mentioned lacks the requirement details.